How many pounds of coffee beans selling for 1.60 per pound should be mixed with 7 pounds of coffee beans selling for 2.40 per pound to obtain a mixture selling for 2.16 per pound
Answer by Theo(13342)

let x = the number of pounds of coffee beans selling for 1.60 per pound.
this makes x + 7 equal to the total number of pounds of coffee beans selling for 2l.16 per pound.
if you add x * 1.6 and you add 7 * 2.40, you will get (x+7) * 2.16
your equation becomes:
1.6x + 2.4*7 = 2.16(x+7)
simplify to get:
1.6x + 16.8 = 2.16x + 15.12
subtract 1.6x from both sides of the equation and subtract 15.12 from both sides of the equation to get:
16.8 - 15.12 = 2.16x - 1.6x
simplify to get:
1.68 = .56x
divide both sides by .56 to get:
1.68/.56 = x
simplify to get:
3 = x which is the same as x = 3.
that's your answer.
3*1.6 + 7*2.4 = 10*2.16
simplify to get:
4.8 + 16.8 = 21.6
simplify further to get:
21.6 = 21.6
answer is confirmed as true.